### v1

# Theoretical study of transition energies and matrix elements of the cadmium atom

Abstract
[…]
I Introduction
Cadmium atom vapor [1, 2, 3, 4, 5] has been selected as one of the nonlinear media for the generation of vacuum ultraviolet (VUV) laser light through four-wave mixing (FWM) method. However, to design and conduct the corresponding experiments, and to accurately estimate the power of the generated light, various atomic properties, such as transition rates between the states involved in the mixing scheme, are required. While these rates have been measured for a wide range of transitions in other alkaline-earth and divalent-like atoms, no experimental or theoretical data are available for transition rates involving high-lying states of neutral Cd [6].
The multireference valence-universal relativistic Fock-space coupled-cluster method (FS-CC) accurately accounts for both electron-correlation and relativistic effects with high accuracy and is powerful for calculating different electronic characteristics of heavy atomic and molecular systems [7]. Accurate ionization potentials and excitation energies have been calculated for various atoms [8, 9, 10, 11, 12, 13, 14, 7](see also references therein). However, to our knowledge, Fock-space coupled cluster method has not been used as extensively for calculating transition matrix elements as it has for transition energies.
[…]
In the present work, we apply the Fock-space coupled cluster method to perform calculation of the transition dipole matrix elements for the $5s^{2}\>{}^{1}S\to 5snp\>^{1,3}P^{o}$, $5s6s\>^{1}S\to 5snp\>^{1,3}P^{o}$ and $5s5d\>^{1}D\to 5snp\>^{1,3}P^{o}$ transitions with $n$ up to 14 for transitions from the $5s^{2}\>{}^{1}S$ and $5s6s\>^{1}S$ states and up to 10 with transitions from the $5s5d\>^{1}D$ state. The calculations are also carried out with the CI+MBPT method. Good agreement between the two methods is found for a set of relatively low-lying transitions with $n$ up to 10, whereas calculations for transitions involving higher-lying states give rather different results. The agreement of the calculated transition energies and matrix elements for the transitions involving low-lying states validates the correctness of the performed computation. The obtained results are expected to be of great value for improving the efficiency of fourth-wave generation in the two FWM schemes previously employed [1, 2, 3, 4, 5].
II Theory and calculation details
II.1 Fock-space coupled cluster
For performing the calculations of the electronic structure of Cd, we used the relativistic version of the Fock-space multireference coupled cluster theory (see, for instance, [22, 23, 24, 7] and references therein). In the present method, the exact multi-electron wave functions are expressed in terms of a wave operator $\Omega$ acting on model vectors $\tilde{\psi}_{i}$:
$|\psi_{i}\rangle=\Omega|\tilde{\psi}_{i}\rangle,$ (1)
where the wave operator takes a form
$\Omega=\bigl{\{}exp(\hat{T})\bigr{\}}.$ (2)
Here, curly brackets denote normal ordering of the operator, and $\hat{T}$ is the so-called cluster operator. This operator takes the form:
$\hat{T}=\sum_{k\leq n_{h}}\sum_{l\leq n_{p}}\hat{T}^{(k,l)},$ (3)
with
$\hat{T}^{(k.l)}=\sum_{\alpha\beta}t^{(k,l)}_{\alpha\beta}a^{\dagger}_{\alpha}a_{\beta},$ (4)
where $t^{(k,l)}_{\alpha\beta}$ is the cluster amplitude associated with the excitation $a^{\dagger}_{\alpha}a_{\beta}$, $a^{\dagger}_{\alpha}$ and $a_{\beta}$ denote sequences of creation and annihilation operators. The quantities $n_{h}$ and $n_{p}$ represent the numbers of holes and particles, respectively, for the target $n_{h}hn_{p}p$ sector of Fock space with respect to the Fermi vacuum. By finding the solution of the equations with respect to the cluster amplitudes, one obtains the corresponding eigenvalues (i.e., energies) and wave functions for the multi-electron problem.
Atomic bispinors were obtained for this ion within a Dirac-Hartree-Fock calculation of the Cd2+ core under the Dirac-Coulomb Hamiltonian. The Fock-Space coupled cluster method with single and double cluster amplitudes (FS-CCSD), within the Dirac-Coulomb Hamiltonian, was employed. We adopted the uncontracted Dyall’s all-electron quadruple-zeta AE4Z basis set [25, 26, 27]. To properly describe higher-lying states, the AE4Z basis set was augmented with 17 $s$-type, 18 $p$-type, 6 $d$-type, and 2 $f$-type diffuse functions as well as 4 $d$-type, and 2 $f$-type core functions. The total basis size was $50s43p27d14f6g3h$. We correlated all electrons of Cd and all the virtual orbitals in the FSCC calculations. The $0h0p$ Fock space sector corresponds to the Cd2+ ion. The wave functions of the interest belong to the $0h2p$ sector. To avoid the typical problem of intruder states, the denominator shifting technique [28] was used. This technique replaces the energy denominators arising in the equation for the cluster amplitudes $t^{(l,k)}_{\alpha\beta}$ with the shifted ones in the following way:
$D_{K}\longrightarrow D_{K}^{{}^{\prime}}(n)=D_{K}+S_{K}\Big{(}\frac{S_{K}}{D_{K}+S_{K}}\Big{)}^{n}.$ (5)
Here $D_{K}$ is the initial energy denominator, $\mathnormal{S_{K}}$ is a shift parameter (in the present case, we resorted to real shifts [28]) and $\mathnormal{n}$ is a non-negative integer parameter. Note that $\mathnormal{S_{K}}$ should be chosen to be relatively small compared to large $\mathnormal{D_{K}}$ values, while still being substantial enough to ensure that the shifted denominators remain negative and sufficiently far from zero to avoid divergences. A very large active space was considered in order to obtain the highly lying excited states, including the $n=(5-9)s$, $n=(5-8)d$, and $n=(5-16)p$ particle orbitals, accounting for 110 active bispinors. The minimum energy-shift parameters that allowed for convergence within this active space were --0.1 and --0.25 for the $0h1p$ and $0h2p$ sectors, respectively. Previously, the denominator shifting technique was successfully developed [29, 30, 31, 32] and employed [33] to avoid divergences for electronic structure calculations in $0h2p$ Fock space sector. For a more detailed and comprehensive description of the technique, one may refer to Ref. [7].
To refine the results of the computation, we also performed calculation within a framework of the Intermediate-Hamiltonian (IH) Fock-space coupled cluster method for incomplete main model spaces [34]. This approach involves partitioning the model space into a main model space and an intermediate model space, such that the contributions of the intermediate space functions to the target states are minimal, thereby avoiding the intruder-state problem. For more details, see Ref. [34]. In this calculation, the total model space consisted of 224 lowest-lying particle orbitals, while the main model space was chosen to comprise determinants with configurations $n=(5-7)s$, $n=(5-7)d$, and $n=(5-14)p$ in sector $0h1p$ and configurations $5sn$, where $n=(5-7),(5-7)d,(5-14)p$, in sector $0h2p$.
The basis set correction for energy calculation was estimated as follows: correlation calculation for the target states was performed using the equation-of-motion scalar-relativistic coupled cluster (EOM-RCC) within the modified AE4Z basis set (comprising of $48s42p27d14f6g3h$ functions) and modified AE4Z basis set with inclusion of even more higher-harmonic’s functions (comprising of $48s42p27d14f8g7h5i$ functions). In this calculation, 2 $s$-type and 1 $p$-type core functions were removed, compared to the FS-CC calculation, to achieve convergence using CFOUR [35] program package. The additional set of $g$,$h$, and $i$ functions was generated using the compact basis set construction technique [36, 37]. In these calculations, the scalar-relativistic variant of the exact two-component (X2C-1e) technique [38] was used. The basis set correction for transition energies was then calculated as a difference between calculation in two above-mentioned basis sets.
For the correlation calculation of transition matrix elements, the framework of the theory of effective operators [39, 40] was employed. This method implies the direct calculation of transition matrix elements, truncating the infinite exponential sum for the wave operator acting on the model space wave functions. In our calculations, we used quadratic truncation level [40, 41]. The calculation of the property integrals was performed using a code developed in Ref. [42]. It carries out the calculation of property integrals over one-electron orbitals obtained in the Dirac-Hartree-Fock (DHF) calculation. The relativistic electron-structure calculations were performed using the dirac [43, 44], Exp-T [45, 46], and CFOUR [35] program packages.
II.2 Configuration interaction
The CI+MBPT method is a hybrid approach combining CI that takes into account an interaction between valence electrons and a method accounting for core-valence correlations using an MBPT operator [47, 48]. In this work, the CI+MBPT calculation is implemented by using the AMBiT code [49]. The first step is to solve the Dirac-Hartree-Fock (DHF) solution for core and valence electrons in the $V^{N}$, $V^{N-1}$, or $V^{N-M}$ approximations, where $N$ is the number of electrons and $M$ is the number of valence electrons. In either choice of potential, the resulting one-electron Dirac-Fock operator is
$h_{DF}=c\vec{\alpha}\cdot\vec{p}+(\beta+1)c^{2}-\frac{Z}{r}+V^{DF},$ (6)
where $\vec{\alpha}$ and $\beta$ are Dirac matrices, and $V^{DF}$ is either of $V^{N}$, $V^{N-1}$, or $V^{N-M}$. The Breit and QED interactions are written into the Dirac-Fock operator. The Breit includes both Gaunt and retardation terms in the frequency-independent limit by
$B_{ij}=-\frac{1}{2r_{ij}}(\vec{\alpha}_{i}\cdot\vec{\alpha}_{j}+(\vec{\alpha}_{i}\cdot\vec{r}_{ij})(\vec{\alpha}_{j}\cdot\vec{r}_{ij})/r^{2}_{ij}).$ (7)
The remaining valence and virtual orbitals are constructed as a linear combination of B-spline basis functions. In the next stage of the AMBiT code, the CI-space of CSFs are formed by allowing an arbitrary number of electron and/or hole excitations from a set of ‘leading configuration’ up to some maximum principal quantum number, $n$, and orbital angular momentum, $l$. The AMBiT code implements an approach, referred as ‘emu CI’ [50], that can reduce the computational cost associated with CI further. In the emu CI approach, in addition to the ‘large-side’ CSFs, there are ‘small-side’ CSFs, formed by allowing electron and/or hole-excitations from a set of leading configurations up to some $n$ and $l$ that can be less that those for large-side CSFs. Finally, the MBPT operator and corrections to energies are implemented in the AMBiT code that can include the core-valence interaction perturbatively.
In the Cd AMBiT calculation, the DHF calculation is carried out with $V^{N-2}$ potentials. The valence basis set is chosen to be $20spdf$ that includes $s$-, $p$-, $d$-, and $f$-orbitals with $n<20$. The $5s^{2}$ and $5s5p$ configurations consists of the set of leading configuration. The single and double excitations from the leading configurations up to $20spdf$ comprise the large-side CSFs of the CI space. The small-side CSFs are constructed by allowing the single and double excitations from the leading configurations up to $6sp5d4f$ in addition to the single excitation from those leading configurations up to $20spdf$. In all AMBiT calculations, the single and double excitations from the $4d$ core are considered and the MBPT calculations are performed with all one-, two-, and three-body MBPT diagrams included based on the $30spdfg$ basis.
III Results and discussion
The calculated transition energies for the considered states are presented in Table 1.
Table 1: Electronic transition energies for the considered states relative to the ground state (in cm-1). Column labeled ‘‘FS-CCSD (shifts)’’ shows the results obtained using the FS-CC method with regular energy-denominator shifts, column labeled "FS-CCSD (IH)" shows the results obtained using the Intermediate-Hamiltonian FS-CC method.
State FS-CCSD (shifts) FS-CCSD (IH) Basis correction Final (FS-CCSD) CI+MBPT Experiment [6]
$5s5p\>^{3}P^{o}_{0}$ 29704 29679 +32 29711 30126 30114
$5s5p\>^{3}P^{o}_{1}$ 30263 30235 +32 30267 30739 30656
$5s5p\>^{3}P^{o}_{2}$ 31472 31435 +32 31467 31862 31827
$5s5p\>^{1}P^{o}_{1}$ 43674 43639 +21 43660 43383 43692
$5s6s\>^{1}S_{0}$ 52756 52725 +81 52806 53806 53310
$5s6p\>^{3}P^{o}_{0}$ 57864 57807 +83 57890 58771 58391
$5s6p\>^{3}P^{o}_{1}$ 57939 57882 +83 57965 58854 58462
$5s6p\>^{3}P^{o}_{2}$ 58132 58070 +83 58153 59027 58636
$5s5d\>^{1}D_{2}$ 58717 58673 +66 58739 59566 59219
$5s6p\>^{1}P^{o}_{1}$ 59489 59445 +76 59521 60499 59907
$5s7p\>^{3}P^{o}_{0}$ 64485 64420 +87 64507 65364 64996
$5s7p\>^{3}P^{o}_{1}$ 64512 64447 +87 64534 65394 65026
$5s7p\>^{3}P^{o}_{2}$ 64584 64516 +87 64603 65490 65094
$5s7p\>^{1}P^{o}_{1}$ 65023 64966 +85 65051 65854 65501
$5s8p\>^{3}P^{o}_{0}$ 67319 67251 +88 67339 68197 67830
$5s8p\>^{3}P^{o}_{1}$ 67332 67264 +88 67352 68212 67842
$5s8p\>^{3}P^{o}_{2}$ 67367 67298 +88 67386 68245 67875
$5s8p\>^{1}P^{o}_{1}$ 67566 67502 +87 67589 68415 68059
$5s9p\>^{3}P^{o}_{0}$ 68812 68743 +89 68832 69684 69314
$5s9p\>^{3}P^{o}_{1}$ 68820 68750 +89 68839 69692 69321
$5s9p\>^{3}P^{o}_{2}$ 68840 68769 +89 68858 69711 69340
$5s9p\>^{1}P^{o}_{1}$ 68947 68880 +88 68968 69801 69439
$5s10p\>^{3}P^{o}_{0}$ 69712 69642 +89 69731 70562 70191
$5s10p\>^{3}P^{o}_{1}$ 69716 69647 +89 69736 70567 70196
$5s10p\>^{3}P^{o}_{2}$ 69729 69659 +89 69748 70579 70208
$5s10p\>^{1}P^{o}_{1}$ 69793 69725 +89 69814 70633 70267
The results of the transition matrix element calculation using the FS-CCSD and CI+MBPT methods, along with their comparison to the existing theoretical and experimental data, are presented in Tables 2 and 3. In Table 2 only the ‘strong’ transitions (transition matrix element value is larger than 0.1 a.u.) are presented.
Table 2: Calculated values of reduced E1 matrix elements (in a.u.) for neutral Cd and comparison to available theoretical and experimental data (the experimental values are deduced from the measured oscillator strengths).
Transition FS-CCSD (shifts) FSCCSD (IH) CI+MBPT Ref. [15] Ref. [16]a Ref. [19]b Ref. [21] Experiment [6]
$5s^{2}\>{}^{1}S_{0}\to 5s5p\>^{3}P^{o}_{1}$ 0.153 0.152 0.169 0.158 0.167 0.14
$5s^{2}\>{}^{1}S_{0}\to 5s5p\>^{1}P^{o}_{1}$ 3.462 3.462 3.552 3.435 3.440 3.4292 3.479 3.01
$5s^{2}\>{}^{1}S_{0}\to 5s6p\>^{1}P^{o}_{1}$ 0.721 0.725 0.697 0.689 0.655 0.670
$5s^{2}\>{}^{1}S_{0}\to 5s7p\>^{1}P^{o}_{1}$ 0.328 0.334 0.319
$5s^{2}\>{}^{1}S_{0}\to 5s8p\>^{1}P^{o}_{1}$ 0.159 0.203 0.186
$5s^{2}\>{}^{1}S_{0}\to 5s9p\>^{1}P^{o}_{1}$ 0.417 0.140 0.126
$5s^{2}\>{}^{1}S_{0}\to 5s10p\>^{1}P^{o}_{1}$ 0.366 0.104 0.094
$5s6s\>^{1}S_{0}\to 5s5p\>^{1}P^{o}_{1}$ 4.145 4.152 3.945 3.923
$5s6s\>^{1}S_{0}\to 5s6p\>^{3}P^{o}_{1}$ 0.815 0.799 0.673
$5s6s\>^{1}S_{0}\to 5s6p\>^{1}P^{o}_{1}$ 7.877 7.887 7.879
$5s6s\>^{1}S_{0}\to 5s7p\>^{1}P^{o}_{1}$ 1.157 1.176 1.152
$5s6s\>^{1}S_{0}\to 5s8p\>^{1}P^{o}_{1}$ 0.512 0.481 0.437
$5s6s\>^{1}S_{0}\to 5s9p\>^{1}P^{o}_{1}$ 0.423 0.269 0.234
$5s6s\>^{1}S_{0}\to 5s10p\>^{1}P^{o}_{1}$ 0.340 0.174 0.149
$5s6s\>^{1}S_{0}\to 5s11p\>^{1}P^{o}_{1}$ 0.160 0.124 0.106
$5s5d\>^{1}D_{2}\to 5s5p\>^{3}P^{o}_{1}$ 0.267 0.269 0.200
$5s5d\>^{1}D_{2}\to 5s5p\>^{1}P^{o}_{1}$ 5.562 5.587 5.353 5.427 6.23
$5s5d\>^{1}D_{2}\to 5s6p\>^{3}P^{o}_{1}$ 1.777 1.761 1.335
$5s5d\>^{1}D_{2}\to 5s6p\>^{1}P^{o}_{1}$ 12.128 12.162 12.296
$5s5d\>^{1}D_{2}\to 5s7p\>^{3}P^{o}_{1}$ 0.313 0.312 0.296
$5s5d\>^{1}D_{2}\to 5s7p\>^{1}P^{o}_{1}$ 0.558 0.563 0.609
$5s5d\>^{1}D_{2}\to 5s8p\>^{1}P^{o}_{1}$ 0.403 0.217 0.232
$5s5d\>^{1}D_{2}\to 5s9p\>^{1}P^{o}_{1}$ 0.016 0.123 0.129
aThe values obtained within CI+All method, where available, are presented.
bThe values obtained using high-order one-body and two-body core-polarization potentials are presented.
As seen in Table 2, the present calculations agree well with the available theoretical and experimental data for the low-lying transitions. For the $5s^{2}\>{}^{1}S_{0}\to 5s5p\>^{3}P^{o}_{1}$ transition, both FS-CCSD approaches give values closer to the experimental one. In the case of $5s^{2}\>{}^{1}S_{0}\to 5s5p\>^{1}P^{o}_{1}$ transition, all theoretical results are quite close, while having a distinct gap with the available experimental value. In contrast, the $5s^{2}\>{}^{1}S_{0}\to 5s6p\>^{1}P^{o}_{1}$ transition values given by both FS-CCSD methods are noticeably far from the very close values given by different versions of the CI method. The largest discrepancy for this transition is found between the IH FS-CCSD and DFCP [19] methods, and amounts to about 10%. For the $5s6s\>^{1}S_{0}\to 5s5p\>^{1}P^{o}_{1}$ transition, the present CI calculation agrees well with the result from Ref. [21]. The difference with the corresponding FS-CCSD calculations is about 5%. The values obtained for the $5s5d\>^{1}D_{2}\to 5s5p\>^{1}P^{o}_{1}$ transition closely match other available theoretical results, with the FS-CCSD calculations being closer to the isolated experimental value.
Table 3: Calculated values of reduced E1 matrix elements (in a.u.) for neutral Cd.
Transition FS-CCSD (shifts) FS-CCSD (IH) CI+MBPT
$5s^{2}\>{}^{1}S_{0}\to$
$5s6p\>^{3}P^{o}_{1}$ 0.005 0.005 0.009
$5s7p\>^{3}P^{o}_{1}$ 0.003 0.002 0.008
$5s8p\>^{3}P^{o}_{1}$ 0.001 0.001 0.005
$5s9p\>^{3}P^{o}_{1}$ 0.033 0.0008 0.003
$5s10p\>^{3}P^{o}_{1}$ 0.023 0.0007 0.003
$5s11p\>^{3}P^{o}_{1}$ 0.012 0.0008 0.003
$5s11p\>^{1}P^{o}_{1}$ 0.089 0.083 0.074
$5s12p\>^{3}P^{o}_{1}$ 0.036 0.0008 0.006
$5s12p\>^{1}P^{o}_{1}$ 0.107 0.129 0.060
$5s13p\>^{3}P^{o}_{1}$ 0.028 0.0003 0.007
$5s13p\>^{1}P^{o}_{1}$ 0.170 0.171 0.051
$5s14p\>^{3}P^{o}_{1}$ 0.001 0.0003 0.022
$5s14p\>^{1}P^{o}_{1}$ 0.031 0.070 0.052
$5s6s\>^{1}S_{0}\to$
$5s5p\>^{3}P^{o}_{1}$ 0.087 0.087 0.091
$5s7p\>^{3}P^{o}_{1}$ 0.029 0.028 0.039
$5s8p\>^{3}P^{o}_{1}$ 0.006 0.005 0.005
$5s9p\>^{3}P^{o}_{1}$ 0.013 0.004 0.002
$5s10p\>^{3}P^{o}_{1}$ 0.010 0.005 0.003
$5s11p\>^{3}P^{o}_{1}$ 0.010 0.005 0.002
$5s12p\>^{3}P^{o}_{1}$ 0.029 0.008 0.002
$5s12p\>^{1}P^{o}_{1}$ 0.026 0.108 0.080
$5s13p\>^{3}P^{o}_{1}$ 0.029 0.009 0.015
$5s13p\>^{1}P^{o}_{1}$ 0.007 0.158 0.063
$5s14p\>^{3}P^{o}_{1}$ 0.008 0.004 0.033
$5s14p\>^{1}P^{o}_{1}$ 0.059 0.216 0.060
$5s5d\>^{1}D_{2}\to$
$5s8p\>^{3}P^{o}_{1}$ 0.165 0.090 0.097
$5s9p\>^{3}P^{o}_{1}$ 0.042 0.003 0.055
$5s10p\>^{3}P^{o}_{1}$ 0.026 0.006 0.038
$5s10p\>^{1}P^{o}_{1}$ 0.008 0.072 0.085
If we analyze the results obtained for the transitions involving higher-lying states from Table 3, we can see that, in general, two different coupled cluster approaches agree to a great extent for transitions from all 3 considered states ($5s^{2}\>{}^{1}S$, $5s6s\>^{1}S$ and $5s5d\>^{1}D$) to the states up to $5s8p\>^{3}P^{o}_{1}$. In the case of transitions from the $5s^{2}\>{}^{1}S_{0}$ state, the correspondence between the values for singlet $\to$ singlet transitions given by the FS-CC method with energy-denominator shifts (denoted as FS-CC (shifts)) and CI method is good up to $5s8p\>^{1}P^{o}_{1}$. For the IH FS-CC method, the agreement with CI extends up to the $5s11p:^{1}P^{o}_{1}$ state. However, for the forbidden singlet $\to$ triplet transitions the situation is much less satisfactory: even for the $5s^{2}\>{}^{1}S_{0}\to 5s6p\>^{3}P^{o}_{1}$ transition the values obtained by the coupled-cluster approaches differ by a factor of two compared to the CI method. Since the transition matrix elements for this transition are already quite small, we suggest that more thorough computations are needed for all higher singlet $\to$ triplet transitions. For the $5s6s\>^{1}S_{0}\to 5snp\>^{1}P^{o}_{1}$ transitions, good agreement is found up to $n=11$ when comparing IH FS-CCSD and CI+MBPT methods, and up to $n=8$ for FS-CCSD (shifts) and CI+MBPT methods. In contrast, for the $5s6s\>^{1}S_{0}\to 5snp\>^{3}P^{o}_{1}$ transitions, the two methods yield noticeably different results for $n>8$. Good agreement between the IH FS-CC and CI methods for singlet $\to$ singlet transitions is also observed for the transitions up to $5s5d\>^{1}D_{2}\to 5s10p\>^{1}P^{o}_{1}$, although in the case of singlet $\to$ triplet transitions, the discrepancy of less than 10% is achieved only for transitions up to the $5s5d\>^{1}D_{2}\to 5s8p\>^{3}P^{o}_{1}$ transition, with a more noticeable difference still present for the $5s5d\>^{1}D_{2}\to 5s5p\>^{3}P^{o}_{1}$.
Comparing the performance of the FS-CC (shifts) and IH FS-CC methods, it can be seen from Table 2 that a relatively good agreement between these implementations for singlet $\to$ singlet transitions is achieved up to the $5s^{2}/5s6s\>^{1}S_{0}\to 5s9p\>^{1}P^{o}_{1}$ transition for those involving the $5s^{2}\>{}^{1}S_{0}$ and $5s6s\>^{1}S_{0}$ states and up to the $5s5d\>^{2}D_{2}\to 5s8p\>^{1}P^{o}_{1}$ for those involving the $5s^{2}\>{}^{1}S_{0}$ state. In these cases, the IH FS-CC method yields results that are generally closer to the CI+MBPT values, particularly for higher-lying transitions. For singlet $\to$ triplet transition, except for the transitions involving low-lying states up to the $5s8p\>^{3}P^{o}_{1}$, all three methods give noticeably different results, with the values computed by IH FS-CC generally showing better agreement with those obtained from CI+MBPT.
In the case of linearly polarized light, the third-order nonlinear susceptibility per atom, $\chi_{a}^{(3)}$, can be expressed as [51]
$\chi_{a}^{(3)}=\frac{1}{6\epsilon_{0}\hbar^{3}}S(\omega_{1}+\omega_{2})\chi_{12}\chi_{34},$ (8)
where
$\chi_{12}=\sum_{l}\left(\frac{\mu_{rl}\mu_{lg}}{\omega_{lg}-\omega_{1}}+\frac{\mu_{rl}\mu_{lg}}{\omega_{lg}-\omega_{2}}\right),$ (9)
$\chi_{34}=\sum_{m}\left(\frac{\mu_{rm}\mu_{mg}}{\omega_{mg}-\omega_{4}}+\frac{\mu_{rm}\mu_{mg}}{\omega_{mg}+\omega_{3}}\right),$ (10)
and $S(\omega_{1}+\omega_{2})$ describes the shape of the two-photon resonance. Here, $\omega_{i}$ is the laser angular frequency, $\mu_{ij}$ are the $z$-components of the transition matrix elements coupling states $i$ and $j$. The summation runs over all intermediate states $l(m)$ (in the considered mixing schemes, these states are $n\>^{1,3}P^{o}_{1}$), whose transition frequencies to the ground state $g$ are denoted as $\omega_{lg}(\omega_{mg})$.
Using the matrix elements calculated in the present work and the same laser frequencies ($\omega_{1}=\omega_{2}=375$ nm, $\omega_{3}=710$ nm), along with the same parameters for the shape factor $S(\omega_{1}+\omega_{2})$ for the two FWM schemes (using $5s6s\>^{1}S_{0}$ or $5s5d\>^{2}D_{2}$ as an intermediate state), considered in Ref. [1] for the spectroscopy of the 229Th isomer, the third-order nonlinear susceptibility $\chi_{a}^{(3)}$ of Cd vapor can be computed. The results obtained in the case of two-photon resonance for the pure 114Cd isotope are presented in Table 4.
Table 4: Absolute values of the third-order nonlinear susceptibility of Cd vapor $|\chi_{a}^{(3)}|$ in units of $10^{-6}(ea_{0})^{4}$cm3 and the corresponding values of factors $|\chi_{12}|$ in units of $10^{-15}(ea_{0})^{2}$z-1 and $|\chi_{34}|$ in units of $10^{-16}(ea_{0})^{2}$z-1 for the pure 114Cd utilizing two FWM schemes with $5s6s\>^{1}S_{0}$ and $5s5d\>^{2}D_{2}$ as intermediate states.
FS-CCSD (shifts) FS-CCSD (IH) CI+MBPT
$5s6s\>^{1}S_{0}$ scheme
$|\chi_{12}|$ 2.32 2.34 2.29
$|\chi_{34}|$ 4.67 6.34 6.29
$|\chi^{(3)}_{a}|$ 1.10 1.50 1.46
$5s5d\>^{2}D_{2}$ scheme
$|\chi_{12}|$ 2.52 2.53 2.48
$|\chi_{34}|$ 6.91 6.57 6.27
$|\chi^{(3)}_{a}|$ 1.59 1.52 1.42
It can be seen that a notable agreement of 5% is achieved for the nonlinear susceptibility $|\chi_{a}^{(3)}|$ and the corresponding factors $|\chi_{12}|$ and $|\chi_{34}|$ obtained using the IH FS-CC and CI+MBPT approaches for both considered FWM schemes. The values given by the FS-CC (shifts) method, however, while showing a 10% agreement for the scheme utilizing the $5s5d\>^{2}D_{2}$ intermediate state, exhibit a discrepancy of about 25% compared to the IH FS-CC and the CI+MBPT methods for the $5s6s\>^{1}S_{0}$ FWM scheme, thus, signifying the advantage of the IH FS-CC technique.
IV Conclusions
We performed calculations of transition dipole matrix elements in neutral Cd for a large set of transitions using the Fock-space coupled cluster and configuration interaction with many-body perturbation theory (CI+MBPT) methods. In the case of Fock-space coupled cluster, two different implementations were used: the denominator shifting technique and the Intermediate Hamiltonian for incomplete model spaces. Previously, similar calculations for cadmium had been carried out only for a few low-lying transitions. Good agreement was found between the utilized methods and available theoretical and experimental results. A correspondence between the results obtained with the FS-CC and CI+MBPT approaches was achieved for a set of $5s^{2}\>{}^{1}S_{0}\to 5snp\>^{1}P^{o}_{1}$ and $5s6s\>^{1}S_{0}\to 5snp\>^{1}P^{o}_{1}$ transitions with $n$ up to 10. In the case of transitions to $5snp\>^{3}P^{o}_{1}$ states, good consistency was observed for transitions with $n<9$. For the transitions involving the $5s5d\>^{1}D_{2}$ state the results obtained with the FS-CC and CI methods are comparable for several low-lying $p$-states with $n<10$.
The values of the third-order nonlinear susceptibility obtained within the IH FS-CCSD and CI+MBPT methods agree within 5% for both considered FWM schemes. As a reference for the $|\chi^{(3)}_{a}|$, we suggest using the IH FS-CCSD results with an estimated uncertainty of 5%. These values are expected to be of significant interest for the preparation of neutral Cd spectroscopy experiment, as well as for the generation of laser light at 148.4 nm through four-wave mixing in cadmium vapor, aimed at high-precision spectroscopy of the 229Th isomeric transition. The discrepancy between the FS-CC and CI results for high-lying transitions highlights the need for more advanced experimental and theoretical approaches to accurately determine considered properties.
Acknowledgments
We thank Qi Xiao and Mikhail Reiter for discussions and comments. Penyazkov and Ding are supported by the National Natural Science Foundation of China (No. 12341401 and No. 12274253) and the Shanghai Municipal Science and Technology Commission (No. 25LZ2600402), Yu is supported by the National Key Research and Development Program of China (2021YFA1402104) and a Project funded by the Space Application System of China Manned Space Program. Skripnikov is supported by Russian Science Foundation under grant  No. 24-12-00092, https://rscf.ru/project/24-12-00092/.